OpenStudy (anonymous):

Points A and B lie on the graph of the equation y = −4.3x. What are the y-coordinates for each point? Point A: (−3, ) Point B: (5, )

OpenStudy (jdoe0001):

hint: \(\large { \begin{array}{llll} A(&-3,\quad )\qquad B(&5,\quad )\\ &\uparrow &\uparrow \\ &{\color{brown}{ x}}=-3&{\color{brown}{ x}}=5 \end{array}\qquad y = −4.3{\color{brown}{ x}} }\)

OpenStudy (anonymous):

okay and...? what else? o.0

OpenStudy (jdoe0001):

for point A, x = -3 so y = -4.3 * x => ? for point B, x = 5, so y = -4.3 * x => ?